OpenStudy (anonymous):

simplify the imaginary number √8

OpenStudy (anonymous):

Do you mean \[-\sqrt{8}\]

OpenStudy (anonymous):

scratch that can you do √-17 please

OpenStudy (anonymous):

Well you can't simplify 17 since it's a prime number

OpenStudy (♪chibiterasu):

You can find the square root of a negative number, it simply doesn't exist.

OpenStudy (anonymous):

cpmplex numbbers they can be simplified

OpenStudy (akashdeepdeb):

\[\sqrt{-17} = \sqrt{17}i\]